# Top Quark Mass Extractions from Energy Correlators: A Feasibility Study

### v1

@@ 1 Introduction
Figure -1019: An illustration of the many subprocesses that enter into the description of a top quark jet substructure observable at the LHC, measured by the detector with varying degrees of angular and energy resolution.
The precise extraction of the top quark mass from a measurement of energy correlators on top quark jets relies on a detailed understanding of the numerous physics subprocesses involved in the production and decay of a boosted top quark, as well as the experimental systematics, as illustrated in figure -1019. To show the feasibility of the measurement, we must show that both the experimental systematics and the dependence on poorly understood nonperturbative subprocess are under control. We do this by systematically investigating the dependence of the proposed measurement on each of the subprocesses in figure -1019 using event generator simulations.
@@ 2 Top Quark Mass Measurements with Energy Correlators
Figure -1018: An illustrative figure produced from Pythia showing the imprint of top and $W$ on the 3-point EEC. First presented in Ref. [27].
In this section, we review our detailed proposal of Ref. [27] to measure the top quark mass in the jet substructure using energy correlators. In short, we propose to extract the top quark mass from a simultaneous fit of two features in the spectrum of the three-point energy correlator. These are shown in figure -1018 as a prominent ridge which imprints the top decay (projected on the front face for a specific value of $\zeta_{S}$) and a second smaller angle peak region formed from the boosted $W$ decay (projected on the back face with the asymmetry cut $\zeta_{A}$ removed). To facilitate these fits, we introduce two projections of the 3-point correlator, $T$ and $W$, each isolating the relevant features of the associated decay.
We define the top projection ($T$) as
$T(\zeta,\zeta_{S},\zeta_{A};m_{W},m_{t},p_{T,{\rm jet}})\equiv$ $\sum_{{}^{\mathrm{hadrons}}_{i,j,k\in{\rm jet}}}\int\textrm{d}\zeta_{ijk}~{}\frac{p_{T,i}\,p_{T,j}\,p_{T,k}}{\big{(}p_{T,\mathrm{jet}}\big{)}^{3}}~{}\frac{\textrm{d}^{3}\sigma_{i,j,k}}{\textrm{d}\zeta_{ijk}}\delta\left(\zeta-\frac{(\sqrt{\zeta_{ij}}+\sqrt{\zeta_{jk}})^{2}}{2}\right)$
@@
$m_{t}=C(\alpha_{s},R)~{}m_{W}\sqrt{\zeta_{t}/\zeta_{W}}+\mathcal{O}\left(\frac{m_{W}}{p_{T,{\rm jet}}},\,\frac{m_{t}}{p_{T,{\rm jet}}}\right),$ (2.4)
@@
Pythia 8.3 1.076 $\pm$ 0.002 1.090 $\pm$ 0.002 1.099 $\pm$ 0.001 1.105 $\pm$ 0.001
Vincia 2.3 1.079 $\pm$ 0.002 1.091 $\pm$ 0.002 1.100 $\pm$ 0.001 1.107 $\pm$ 0.002
Herwig 7.3 Dipole 1.071 $\pm$ 0.002 1.082 $\pm$ 0.001 1.098 $\pm$ 0.001 1.100 $\pm$ 0.002
Herwig 7.3 A.O. 1.094 $\pm$ 0.001 1.106 $\pm$ 0.001 1.112 $\pm$ 0.001 1.125 $\pm$ 0.001
@@ 2.1 Strategy for Peak Finding with Event Generator Data
Without further input from analytical calculations, $\zeta_{t}$ and $\zeta_{W}$ must be found from fits of the peak regions in the $T$ and $W$ distributions. All fits in this paper were performed using fifth-degree polynomials for $\zeta_{t}$ and sixth-degree for $\zeta_{W}$ restricted to the region of half peak maximum. Polynomials of degrees 2 to 10 were studied, and no statistically significant shifts in the extracted peak position were observed at degrees higher than those chosen. Included within the uncertainty on each fit is an additional uncertainty for the ambiguity in the fitting procedure, found from the standard deviation of the extracted peak positions over this stable window of the polynomial degree and the standard deviation due to variation of the fitting window in the spectrum by ±$\pm 10\%$. Nevertheless, these additional fitting uncertainties are always substantially smaller than the statistical uncertainties from the distributions. Other fitting procedures were explored, such as Gaussian, crystal ball, and Breit-Wigner functions, and while consistent results were found, simple polynomial fits proved to be the most numerically stable.
@@ 3 Experimental Reconstruction Uncertainties
In this section, we analyze the effects of uncertainties characterizing the experimental environment on our energy-correlator-based top mass extraction. Throughout this paper we use the Herwig 7.3 [33, 34, 35, 36] [^2] In the first version of Ref. [27], we had employed Herwig 7.2 compared to version 7.3 in this manuscript. We found a markedly improved stability of Herwig 7.3 compared to the previous version. This is reflected in the value of the perturbative constant $C$ in table 1 (which for $R=1.2$, our default choice, is in significantly better agreement with Pythia 8.3 and Vincia 2.3) as well as in the smallness of hadronization corrections displayed in figure -1006 below. For the sake of completeness, a more detailed comparison of versions 7.2 and 7.3 will be presented in the next revision of Ref. [27]., Pythia 8.3 [37] and Vincia 2.3 [38] parton shower simulations, as well as the FastJet [39] implementation of the anti-$k_{T}$ algorithm [40]. We focus on uncertainties in the calibration of the jet energy scale, which affects the reconstructed $p_{T,\mathrm{jet}}$, uncertainties in the constituent energy scale, which shift the momenta of the jet constituents, and different track reconstruction efficiency models that account for a $p_{T}$-dependent loss of particle tracks. The effects of the various uncertainties are estimated by repeating the measurement of the ratio $\zeta_{t}/\zeta_{W}$ while changing the parameter characterizing an uncertainty source by one standard deviation. The result is compared to the nominal analysis in each bin of $p_{T,\mathrm{jet}}$. As a measure of the effect, we calculate the weighted average of the shifts in the extracted top quark mass for all $p_{T,\mathrm{jet}}$ bins.
@@ 3.1 Jet Energy Scale
Figure -1017: Model of the relative CMS jet energy scale uncertainty as a function of the jet $p_{T}$ (left). Average shift in the top quark mass due to the jet energy scale uncertainty (right). The shift is less than 100 MeV.
@@
Figure -1016: The shift in the extracted top quark mass due to changes of the energy scale of charged jet constituents (upper left), photons (upper right), and other neutral constituents (lower). The shift is found to be less than 20 MeV.
To test the presented measurement strategy against realistic variations, we model the CMS jet energy scale uncertainty [46, 47] with the function displayed in figure -1017 (left). Thus, the energy scale of most jets in this study is varied by 0.5–2.5%. We repeat the analysis while changing the jet momenta up and down according to the uncertainty defined above. For this uncertainty, it was chosen to only change the jet momentum without altering its constituents, which are studied separately below. The results are shown in figure -1017 (right). Because of the ratio of the $\zeta_{t}$ and $\zeta_{W}$, the measurement is nearly unaffected by changes in the jet energy scale, which show a shift in the top quark mass below 100 MeV.
@@ 3.2 Constituent Energy Scale
In addition to uncertainties in the energy scale of the jet, we study the effect of varying the constituent momenta. For this, we follow the strategy of an existing measurement using energy correlators inside jets by the CMS Collaboration [21] and separately vary the energy scale by 1% for charged jet constituents, 3% for photons, and 5% for other neutral particles. The results are displayed in figure -1016 and show negligible impact.
Figure -1015: The shift in the extracted top quark mass for two scenarios of track reconstruction inefficiencies for an analysis using all particles (left) and only tracks (right). In all cases, the observed shift is less than 150 MeV. The horizontal dashed line shows the central value of the mass extracted from an inclusive final state at the parton level. A remarkably small shift from the central value is observed when the measurement is restricted to only tracks.
@@ 3.3 Tracking Efficiency
We now focus on the feasibility of performing the proposed top mass measurement on charged particles (tracks) only. This is a unique advantage of the energy correlator measurement since other approaches to the top mass extraction from jet substructure suffer large corrections from the restriction to charged particle final states. In contrast, the restriction to tracks is a small effect to the energy correlator spectrum which can be systematically included in theoretical calculations [48, 49, 50, 51, 52, 53, 54]. This behavior is validated by figure -1015, where the red circles in the right panel show the bin-by-bin top mass extraction performed on tracks. A remarkably small shift ($\lesssim 1~{}$GeV) relative to the central value of the all-hadrons result (the dashed line) is observed. This shift is computable to high precision from analytical theory, and it is exceedingly promising that only $10\%$ accuracy in the description of the nonperturbative track function input would be needed for this shift to be considered negligible.
@@
Figure -1014: The shift in the extracted top quark mass assuming two scenarios of track reconstruction applied only to particles from heavy quark decays or light quark decays. Results are compared for an analysis using all particles (left column) and only tracks (right column) as well as the models ‘SMP-22-015’ (upper row) and ‘EXO-23-002’ (lower row). The horizontal dashed line shows the central value of the mass extracted from an inclusive final state at the parton level. A remarkably small shift from the central value is observed when the measurement is restricted to charged particles.
@@
While track-finding algorithms have a very high efficiency in current experiments, they can deteriorate in dense environments such as highly energetic jets. In this context, we investigate two models, following the uncertainty estimations of two CMS analyses [21, 55]. In model ‘SMP-22-015’ [21] 3% of all tracks within a jet are randomly excluded. For model ‘EXO-23-002’ [55] a $p_{T}$ dependent approach is chosen such that 3% of tracks with $p_{\text{T}}<20\;\text{GeV}$ and 1% of tracks with $p_{\text{T}}>20\;\text{GeV}$ are randomly excluded. The results are shown in figure -1015 for a version of the analysis with all particles (left) and a version using only tracks (right). The observed shift due to different tracking efficiency models is small and consistent with 0 MeV within statistical uncertainties.
@@
A known effect in detectors is the different jet response depending on the origin of a jet. This difference in jet response between jets initiated by gluons, light quarks, and bottom quarks poses a serious challenge to many top quark mass measurements. Because of this, calibrations of the jets used to reconstruct the top quark via the known $W$ boson mass peak suffer from large uncertainties. Since in the measurement strategy discussed in this paper, the distribution associated with the $W$ is used comparably, we test the effect by applying the models ‘SMP-22-015’ and ‘EXO-23-002’ separately to particles that originate from a heavy flavor bottom quark or those that originate from a light flavor quark. The results of the $m_{t}$ extraction are displayed in figure -1014 for an analysis with all particles (left) and a version using only tracks (right). The effect remains small, particularly when the measurement is performed using only tracks.
@@ 4 Sensitivity to subprocesses
Having illustrated that our proposed measurement is minimally affected by experimental reconstruction uncertainties, it remains to be demonstrated that it is under theoretical control in the complicated hadron collider environment. The complete description of top quark production and decay at a hadron collider involves many subprocesses, as illustrated in figure -1019. From the theoretical perspective, a number of these subprocesses, such as color reconnection and the underlying event, are non-universal power corrections to factorization theorems. They are poorly understood and often require input from models. Therefore, to motivate the precision theoretical calculation of energy correlator observables on top quark decays in hadron colliders, it is desirable to show that these poorly understood nonperturbative contributions have a minimal effect on the energy correlator distribution. Furthermore, by studying the dependence of the energy correlator observable to these subprocesses, we can determine where theoretical effort must be concentrated to develop a precision first principles understanding of the proposed measurement.
@@
Figure -1013: The impact of the modelling of color reconnection on the extracted top mass is found to be less than 150 MeV.
@@ 4.1.1 Color Reconnection Modelling
Sensitivity to the color reconnection model, compared to the default, is shown in figure -1013. Variations of the order of $100$ MeV show that our observable is indeed robust to such effects.
@@ 4.1.2 Underlying Event Tune
Figure -1012: The impacts of variations in the underlying event tune on the extracted top mass are found to be less than 100 MeV.
@@
The results of the variation of the underlying event tune are shown in figure -1012 and are found to impact the top quark mass extraction from energy correlators by less than 200 MeV.
@@ 4.2 Sensitivity to Systematically Improvable Subprocesses
Figure -1011: Impact of variations of the PDFs are found to be minimal, showing that the observable isolates properties of the top quark decay.
Figure -1010: Impact of variations of the ISR (left) and FSR (right) scales.

### v2

@@ 1 Introduction
Figure -959: An illustration of the many subprocesses that enter into the description of a top quark jet substructure observable at the LHC, measured by the detector with varying degrees of angular and energy resolution.
The precise extraction of the top quark mass from a measurement of energy correlators on top quark jets relies on a detailed understanding of the numerous physics subprocesses involved in the production and decay of a boosted top quark, as well as the experimental systematics, as illustrated in figure -959. To show the feasibility of the measurement, we must show that both the experimental systematics and the dependence on poorly understood nonperturbative subprocess are under control. We do this by systematically investigating the dependence of the proposed measurement on each of the subprocesses in figure -959 using event generator simulations.
@@ 2 Top Quark Mass Measurements with Energy Correlators
Figure -958: An illustrative figure produced from Pythia showing the imprint of top and $W$ on the 3-point EEC. First presented in Ref. [27].
In this section, we review our detailed proposal of Ref. [27] to measure the top quark mass in the jet substructure using energy correlators. In short, we propose to extract the top quark mass from a simultaneous fit of two features in the spectrum of the three-point energy correlator. These are shown in figure -958 as a prominent ridge which imprints the top decay (projected on the front face for a specific value of $\zeta_{S}$) and a second smaller angle peak region formed from the boosted $W$ decay (projected on the back face with the asymmetry cut $\zeta_{A}$ removed). To facilitate these fits, we introduce two projections of the 3-point correlator, $T$ and $W$, each isolating the relevant features of the associated decay.
We define the top projection ($T$)111 Regarding the experimental measurement, we note that the distribution $T$ (as well as $W$ defined below) and the hadronic cross-sections in the all present formulae are differential in $p_{T,{\rm jet}}$ and should be integrated over the narrow $p_{T,{\rm jet}}$ bins in which the measurements are performed. The differentials have been omitted for brevity. as
$T(\zeta,\zeta_{S},\zeta_{A};m_{W},m_{t},p_{T,{\rm jet}})\equiv$ $\sum_{{}^{\mathrm{hadrons}}_{i,j,k\in{\rm jet}}}\int\textrm{d}\zeta_{ijk}~{}\frac{\textrm{d}\sigma_{i,j,k}}{\textrm{d}\zeta_{ijk}}~{}\frac{p_{T,i}\,p_{T,j}\,p_{T,k}}{\big{(}p_{T,\mathrm{jet}}\big{)}^{3}}~{}\delta\left(\zeta-\frac{(\sqrt{\zeta_{ij}}+\sqrt{\zeta_{jk}})^{2}}{2}\right)$
@@
$m_{t}=m_{W}\left[C(\alpha_{s},R)\sqrt{\zeta_{t}/\zeta_{W}}+\mathcal{O}\left(\frac{m_{W}}{p_{T,{\rm jet}}},\frac{m_{t}}{p_{T,{\rm jet}}}\right)\right],$ (2.4)
@@
Pythia 8.3 1.075 $\pm$ 0.001 1.090 $\pm$ 0.001 1.099 $\pm$ 0.001 1.105 $\pm$ 0.001
Vincia 2.3 1.078 $\pm$ 0.001 1.091 $\pm$ 0.002 1.101 $\pm$ 0.001 1.107 $\pm$ 0.001
Herwig 7.3 Dipole 1.078 $\pm$ 0.001 1.088 $\pm$ 0.001 1.098 $\pm$ 0.001 1.106 $\pm$ 0.001
Herwig 7.3 A.O. 1.092 $\pm$ 0.001 1.104 $\pm$ 0.001 1.113 $\pm$ 0.001 1.120 $\pm$ 0.001
@@ 2.1 Strategy for Peak Finding with Event Generator Data
Without further input from analytical calculations, $\zeta_{t}$ and $\zeta_{W}$ must be found from fits of the peak regions in the $T$ and $W$ distributions. All fits in this paper were performed using fifth-degree polynomials for $\zeta_{t}$ and sixth-degree for $\zeta_{W}$. Sufficiently high number of Monte Carlo events were generated such that the statistical error is subdominant in the analysis below. Several fitting procedures were explored, such as Gaussian, crystal ball, and Breit-Wigner functions, and while consistent results were found, simple polynomial fits proved to be the most numerically stable. For the case of the $W$ distribution more care is required. While the hadron-level $W$ distributions displayed a smooth peaked feature for a range of $p_{T,\rm jet}$ bins, the parton level distributions displayed a small oscillating skew in a localised region around the peak. As a result, the parton-level peak position retains some sensitivity to the choice of degree of polynomial and the fit range. To this end, we restricted to a region roughly defined by half peak maximum and avoided zooming too narrowly around the peak. We discarded polynomials up to degree-five for $W$ distributions and found that degree-six and degree-seven polynomials resulted in equivalent fits. To account for ambiguity associated in the fitting procedure, we incorporate in our uncertainty the differences in the peak positions found via degree-six and degree-seven polynomials, and due to variation of the fitting window in the spectrum by ±$\pm 10\%$. Examples of the polynomial fits and a break down of the errors can be seen in Fig. -957.
Figure -957: Representative fits for peak locations in the $W$ distribution, shown here for Pythia 8.3, $p_{T}\in[500,525]$ GeV for hadron and parton level. The inset shows the uncertainty breakdown, with the dominant uncertainty arising from the variations in the polynomial order.
@@ 3 Experimental Reconstruction Uncertainties
In this section, we analyze the effects of uncertainties characterizing the experimental environment on our energy-correlator-based top mass extraction. Throughout this paper we use the Herwig 7.3 [33, 34, 35, 36] [^3] In the first version of Ref. [27], we had employed Herwig 7.2 compared to version 7.3 in this manuscript. We found a markedly improved stability of Herwig 7.3 compared to the previous version. This is reflected in the value of the perturbative constant $C$ in table 1 (which for $R=1.2$, our default choice, is in significantly better agreement with Pythia 8.3 and Vincia 2.3) as well as in the smallness of hadronization corrections displayed in figure -945 below. For the sake of completeness, a more detailed comparison of versions 7.2 and 7.3 will be presented in the next revision of Ref. [27]., Pythia 8.3 [37] and Vincia 2.3 [38] parton shower simulations, as well as the FastJet [39] implementation of the anti-$k_{T}$ algorithm [40]. We focus on uncertainties in the calibration of the jet energy scale, which affects the reconstructed $p_{T,\mathrm{jet}}$, uncertainties in the constituent energy scale, which shift the momenta of the jet constituents, and different track reconstruction efficiency models that account for a $p_{T}$-dependent loss of particle tracks. The effects of the various uncertainties are estimated by repeating the measurement of the ratio $\zeta_{t}/\zeta_{W}$ while changing the parameter characterizing an uncertainty source by one standard deviation. The result is compared to the nominal analysis in each bin of $p_{T,\mathrm{jet}}$. As a measure of the effect, we calculate the weighted average of the shifts in the extracted top quark mass for all $p_{T,\mathrm{jet}}$ bins.
@@ 3.1 Jet Energy Scale
Figure -956: Model of the relative CMS jet energy scale uncertainty as a function of the jet $p_{T}$ (left). Average shift in the top quark mass due to the jet energy scale uncertainty (right). The shift is less than 100 MeV.
@@
Figure -955: The shift in the extracted top quark mass due to changes of the energy scale of charged jet constituents (upper left), photons (upper right), and other neutral constituents (lower). The shift is found to be less than 100 MeV.
To test the presented measurement strategy against realistic variations, we model the CMS jet energy scale uncertainty [46, 47] with the function displayed in figure -956 (left). Thus, the energy scale of most jets in this study is varied by 0.5–2.5%. We repeat the analysis while changing the jet momenta up and down according to the uncertainty defined above. For this uncertainty, it was chosen to only change the jet momentum without altering its constituents, which are studied separately below. The results are shown in figure -956 (right). Because of the ratio of the $\zeta_{t}$ and $\zeta_{W}$, the measurement is nearly unaffected by changes in the jet energy scale, which show a shift in the top quark mass below 100 MeV.
@@ 3.2 Constituent Energy Scale
In addition to uncertainties in the energy scale of the jet, we study the effect of varying the constituent momenta. For this, we follow the strategy of an existing measurement using energy correlators inside jets by the CMS Collaboration [21] and separately vary the energy scale by 1% for charged jet constituents, 3% for photons, and 5% for other neutral particles. The results are displayed in figure -955 and show negligible impact.
Figure -954: The shift in the extracted top quark mass for two scenarios of track reconstruction inefficiencies for an analysis using all particles (left) and only tracks (right). In all cases, the observed shift is $\sim 200$ MeV. The horizontal dashed line shows the central value of the mass extracted at the parton level.
@@ 3.3 Tracking Efficiency
We now focus on the feasibility of performing the proposed top mass measurement on charged particles (tracks) only. This is a unique advantage of the energy correlator measurement since other approaches to the top mass extraction from jet substructure suffer large corrections from the restriction to charged particle final states. In contrast, the restriction to tracks is a small effect to the energy correlator spectrum which can be systematically included in theoretical calculations [48, 49, 50, 51, 52, 53, 54]. This behavior is validated by figure -954, where the red circles in the right panel show the bin-by-bin top mass extraction performed on tracks. A remarkably small shift ($\lesssim 1~{}$GeV) relative to the central value of the all-hadrons result (the dashed line) is observed. This shift is computable to high precision from analytical theory, and it is exceedingly promising that only $10\%$ accuracy in the description of the nonperturbative track function input would be needed for this shift to be considered negligible.
@@
Figure -953: The shift in the extracted top quark mass assuming two scenarios of track reconstruction applied only to particles from heavy quark decays or light quark decays. Results are compared for an analysis using all particles (left column) and only tracks (right column) as well as the models ‘SMP-22-015’ (upper row) and ‘EXO-23-002’ (lower row). The horizontal dashed line shows the central value of the mass extracted from an inclusive final state at the parton level. A remarkably small shift from the central value is observed when the measurement is restricted to charged particles.
@@
While track-finding algorithms have a very high efficiency in current experiments, they can deteriorate in dense environments such as highly energetic jets. In this context, we investigate two models, following the uncertainty estimations of two CMS analyses [21, 55]. In model ‘SMP-22-015’ [21] 3% of all tracks within a jet are randomly excluded. For model ‘EXO-23-002’ [55] a $p_{T}$ dependent approach is chosen such that 3% of tracks with $p_{\text{T}}<20\;\text{GeV}$ and 1% of tracks with $p_{\text{T}}>20\;\text{GeV}$ are randomly excluded. The results are shown in figure -954 for a version of the analysis with all particles (left) and a version using only tracks (right). The observed shift due to different tracking efficiency models is small.
@@
A known effect in detectors is the different jet response depending on the origin of a jet. This difference in jet response between jets initiated by gluons, light quarks, and bottom quarks poses a serious challenge to many top quark mass measurements. Because of this, calibrations of the jets used to reconstruct the top quark via the known $W$ boson mass peak suffer from large uncertainties. Since in the measurement strategy discussed in this paper, the distribution associated with the $W$ is used comparably, we test the effect by applying the models ‘SMP-22-015’ and ‘EXO-23-002’ separately to particles that originate from a heavy flavor bottom quark or those that originate from a light flavor quark. The results of the $m_{t}$ extraction are displayed in figure -953 for an analysis with all particles (left) and a version using only tracks (right). The effect remains small, particularly when the measurement is performed using only tracks.
@@ 4 Sensitivity to subprocesses
Having illustrated that our proposed measurement is minimally affected by experimental reconstruction uncertainties, it remains to be demonstrated that it is under theoretical control in the complicated hadron collider environment. The complete description of top quark production and decay at a hadron collider involves many subprocesses, as illustrated in figure -959. From the theoretical perspective, a number of these subprocesses, such as color reconnection and the underlying event, are non-universal power corrections to factorization theorems. They are poorly understood and often require input from models. Therefore, to motivate the precision theoretical calculation of energy correlator observables on top quark decays in hadron colliders, it is desirable to show that these poorly understood nonperturbative contributions have a minimal effect on the energy correlator distribution. Furthermore, by studying the dependence of the energy correlator observable to these subprocesses, we can determine where theoretical effort must be concentrated to develop a precision first principles understanding of the proposed measurement.
@@
Figure -952: The impact of the modelling of color reconnection on the extracted top mass is found to be less than 150 MeV.
@@ 4.1.1 Color Reconnection Modelling
Sensitivity to the color reconnection model, compared to the default, is shown in figure -952. Variations of the order of $100$ MeV show that our observable is indeed robust to such effects.
@@ 4.1.2 Underlying Event Tune
Figure -951: The impacts of variations in the underlying event tune on the extracted top mass are found to be less than 100 MeV.
@@
The results of the variation of the underlying event tune are shown in figure -951 and are found to impact the top quark mass extraction from energy correlators by less than 200 MeV.
@@ 4.2 Sensitivity to Systematically Improvable Subprocesses
Figure -950: Impact of variations of the PDFs are found to be minimal, showing that the observable isolates properties of the top quark decay.
Figure -949: Impact of variations of the ISR (left) and FSR (right) scales.